As a fighter pilot makes a tight turn (with his head toward the centre of curvature), the blood drains from his head, leading to loss of brain function. Once the radial acceleration reaches about 4g, the pilot’s vision switches to black and white and narrows to ‘tunnel vision’. If this acceleration is maintained, or increased, the pilot will lose all vision and will black out. A F-22 Raptor pilot in a dogfight throws his aircraft into a tight turn (r = 5.80 km) while flying at Mach 2.0. Is he likely to survive?

Grade:12th pass

1 Answers

NIKHIL CHOTRANI
5 years ago
no he will not survive
the centripetal acceleration must not exceed 4g or 4*10N or 40(m/s2)
nd we know centripetal accn is (v)square/r=(686)square/5800=686*686/5800=81.13(m/s2)
v=686m/s becoz it is mach2.0 means it travels 2 times as faster as sound and speed of sound is 343m/s
and r=5.8 km=5800 m
so here centripetal accn or accn on man is exceeding the safe limit